An electromagnetic wave of intensity 240 W/m2 is incident normally on a rectangular card with sides of 19 cm and 31 cm that reflects all the radiation incident on it. Find the force exerted by the electromagnetic wave on the card if the radiation is incident at an angle of 30
Answer #1

Using the definition of pressure, express the force exerted on the card

Fr = Pr* A*cos(theta)

If the card reflects all of the radiation incident on it, the force exerted on the card is doubled:

so Fr =2* Pr* A*cos(theta)

Pr= I/c

where c= 3*10^8m/sec

Pr= 240/3*10^8

Fr =2* Pr* A*cos(theta)

Fr= 2* (240/3*10^8) *0.19*0.31*cos30

Fr = 8.1614*10^-8N = 81.6nN